OpenStudy (anonymous):

Solve: 2(x + 1) = −8 5 −3 −5 −4.5

OpenStudy (austinl):

Ok, We would first want to distribute on the left side of the equation. Could you do that for me?

OpenStudy (anonymous):

2(-5+1) = -8 x = -5

OpenStudy (austinl):

@NobodySpecial We don't just give answers, nor do we not show our work.

OpenStudy (goformit100):

2(x + 1) = −8 or, 2x + 2 = -8 or, 2x = -10 or, x = -5........................................Ans.

OpenStudy (anonymous):

2(-5 + 1) = -8 is my work....

OpenStudy (austinl):

That isn't work, that is using a guess and check technique. @goformit100 used an algebraic method that I had planned on guiding the asker through.

OpenStudy (anonymous):

Guess and check works fine with multiple choice.

OpenStudy (austinl):

One last thing before I go, Guess and check may work fine for multiple choice. However, to apply this to later problems that do not offer choices for answers; you need to know how to solve the problems algebraically.

OpenStudy (anonymous):

@austinL I agree completely.

OpenStudy (goformit100):

I am Sorry Sir. Please Excuse me for my deeds :'(

OpenStudy (austinl):

@goformit100 You aren't the one that I was referring to...